Please sir explain this question, as I am completely dependent on askiitian. plzzz.......plzz......plz...........sir clear my doubt.A narrow uniform glass tube 80cm long and open at both ends is half immersed in mercury .Then top of the tube is closed and it is taken out of mercury. A column of mercury 22cm long then remains in the tube.What is the atmospheric pressure?(please sir explain how to take pressure in both the conditions. Please sir explain in detail.)

Grade:12th pass

1 Answers

Arun
5 years ago
When the glass tube was in mercury with half dipped Pressure on top of mecury should be equal to mercury pressure Therefore that pressure will be P(1) =13.6g(44)Now after taking tube out thumb is placed on the top so This should apply there also there now in this condition the pressure of mercury + P1= atmospheric pressure applying now on the other open end ThereforeAtmospheric pressure=P1+13.6g(22) = 13.6g(44)+13.6g(22) =13.6g(64)Atmospheric pressure =64 cm of Hg column